OpenStudy (anonymous):

algebra 1 help plz will medal

OpenStudy (anonymous):

when \[9^{\frac{ 2 }{ 3 }}\] is written in simplest radical form, what value remains under the radical

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

we must put this as a improper fraction 3(9)+2 =29 so, 29/3=9.66666